# Semantic Context Matters: Improving Conditioning for Autoregressive Models

### v1

@@ Abstract
Recently, autoregressive (AR) models have shown strong potential in image generation, offering better scalability and easier integration with unified multi-modal systems compared to diffusion-based methods. However, extending AR models to general image editing remains challenging due to weak and inefficient conditioning, often leading to poor instruction adherence and visual artifacts. To address this, we propose SCAR, a Semantic-Context-driven method for AutoregRessive models. SCAR introduces two key components: Compressed Semantic Prefilling, which encodes high-level semantics into a compact and efficient prefix, and Semantic Alignment Guidance, which aligns the last visual hidden states with target semantics during autoregressive decoding to enhance instruction fidelity. Unlike decoding-stage injection methods, SCAR builds upon the flexibility and generality of vector-quantized-based prefilling while overcoming its semantic limitations and high cost. It generalizes across both next-token and next-set AR paradigms with minimal architectural changes. SCAR achieves superior visual fidelity and semantic alignment on both instruction editing and controllable generation benchmarks, outperforming prior AR-based methods while maintaining controllability. All code will be released.
@@ 1 Introduction
The advent of large-scale generative models [sdxl, lan2025flux, he2025ragsr, han2025infinity, liu2025infinitystar] has revolutionized the field of image editing [sdedit, controlnet], opening up unprecedented possibilities for creative expression and content manipulation. Within this rapidly evolving landscape, two dominant generative paradigms have emerged: diffusion [ddpm, dit] and autoregressive (AR) models [llamagen, var]. Diffusion models, renowned for their powerful text-to-image synthesis capabilities, demonstrate a strong grasp of semantic concepts, allowing for flexible and diverse edits based on user prompts. Concurrently, autoregressive models, owing to their excellent scaling properties and immense potential within unified multimodal generation and understanding model (UMM) architectures [lumina-mgpt, zhuang2025vargpt], have achieved generation quality approaching that of top-tier diffusion models, establishing them as a critical and highly promising direction for future research.
@@
However, translating this raw generative power into versatile, general image editing capabilities remains a significant challenge for AR models. Compared to diffusion-based editing methods [instructpix2pix, controlnet, controlnet++, unicontrol], recent AR-based approaches [li2024controlar, mu2025editar] have yet to achieve the same level of semantic control and general editing scope. These works can be broadly categorized into two types based on their conditioning strategies: (1) Decoding-stage injection (Figure 1.a), as in ControlAR [li2024controlar], injects condition signals into the model’s intermediate layers. It performs excellent controllable generation, but the injected strong spatial guidance disrupts the autoregressive process for general editing (see Figure 6). (2) Prefilling-stage conditioning (Figure 1.b), employed by works like EditAR [mu2025editar] and most UMMs  [xin2025lumina, zhuang2025vargpt, li2025onecat, cao2025hunyuanimage], prepends visual tokens from the condition image to the input sequence. However, this method introduces debilitating computational overhead: prepending vector-quantized (VQ) tokens of the condition image can double the sequence length, increasing the computational cost, as shown in Figure 1.c. Moreover, VQ tokens are widely recognized as sparse-semantic [harmon, blip3, unitok], lacking the high-level representation needed for complex edits.
@@
First, we propose Compressed Semantic Prefilling, which replaces the inefficient VQ token prefix with compact, rich-semantic vision features extracted from the source image via a frozen vision foundation model (VFM) [dinov2, clip, vit]. We introduce a learnable semantic compression module that drastically reduces the prefilling sequence length (e.g., $4\times$ compression from 1024 to 256 tokens) while preserving essential high-level semantics for editing. This significantly enhances computational efficiency and provides the model with a compact semantic understanding of the source content before generation, as shown in Figure 1.d. Notably, our semantic prefix remains robust even when compressed, while VQ token prefix suffers sharp performance drops under similar settings.
Second, to bridge the gap between sparse textual instructions and the desired dense visual guidance, we introduce a novel Semantic Alignment Guidance. Instead of relying solely on the sparse text condition, we use the visual features of the target edited image from the VFM as a dense semantic guide. Unlike the prior method [mu2025editar] that distills supervision onto output VQ tokens, we apply an auxiliary constraint objective that aligns the autoregressive model’s last visual hidden states with the target’s semantic representation, providing a formulation more consistent with the causal decoding process. This provides a dense, in-context learning signal that steers the model’s internal reasoning toward the edit target.
@@ 2.1 Image Generation
Two dominant paradigms have emerged in image generation: diffusion models and AR models. Diffusion models synthesize images by iteratively denoising Gaussian noise, with DDPM [ho2020denoising] marking a major breakthrough. Subsequent work improves generation quality and efficiency [nichol2021improved, rombach2022high, zhang2025robust] through advances in sampling strategies and latent-space modeling. They have become the backbone of text-to-image and text-to-video generation [saharia2022photorealistic, singer2022make, ma2024passersby, lan2025flux, zhang2025boow], typically using U-Net for denoising and CLIP [clip, wei2025hq] or T5 [t5] for text conditioning via cross-attention. Recent models like DiT [dit] replace U-Net with Transformer backbones, achieving strong performance. Despite their success, diffusion models remain computationally expensive, motivating exploration of more efficient alternatives such as autoregressive approaches.
@@
In contrast, autoregressive (AR) models formulate image synthesis as a sequence modeling problem, predicting visual tokens step by step. Early works [van2016pixel] focused on pixel-level generation. Building on the success of large language models [touvron2023llama, achiam2023gpt], recent approaches adopt discrete quantizers such as VQ-VAE [van2017neural] and VQ-GAN [esser2021taming] to convert image patches into token indices, enabling next-token prediction over visual sequences. Recent advances have expanded this paradigm in two directions: (1) conventional next-token AR models [open-magvit2, ma2025towards, ma2025stage], such as LlamaGen [llamagen] and AiM [aim], leverage large-scale transformer backbones for high-quality image synthesis; and (2) Visual Autoregressive Modeling (VAR), which reformulates generation as next-set prediction rather than raster-scan token prediction, offering improved scalability and global coherence [var, ma2024star, tang2024hart, han2025infinity, zhuang2025vargpt]. Overall, AR-based approaches have achieved image quality comparable to diffusion models while being more efficient in sampling and deployment.
@@ 2.2 General Image Editing
General Image editing, including controllable and instruction editing, is a key focus in generative modeling, guided by external signals (e.g., text, edge maps, reference images). While early approaches incorporated class labels or attributes in GANs [mirza2014conditional, song2023fashion] and VAEs [kingma2013auto, zeng2024cat], recent diffusion-based methods [controlnet, t2i-adapter, unicontrol] achieve fine-grained control via cross-attention or adapter modules. Building on this, text-guided instruction editing methods [prompt2prompt, sdedit, instructpix2pix] further extend controllable generation to image manipulation. However, these approaches often involve costly inversion, require modality-specific tuning, or suffer from limited generalization across tasks and conditions.
In contrast, image editing with autoregressive (AR) models, including next-token and next-set paradigms, remains relatively underexplored. Recent AR-based methods can be broadly categorized by the stage at which control signals are injected: Decoding-stage injection [yao2024car, li2024controlar, xu2025scalar] introduces control signals dynamically during generation, typically via cross-attention or feature modulation. These approaches offer strong controllability for pixel-level generation, but their complex architectural modifications often limit generalization to instruction-driven editing. Prefilling-stage conditioning [lumina-mgpt, xin2025lumina, zhuang2025vargpt, mu2025editar, qu2025varsr, controlvar] prepends visual tokens from the source image to the input sequence, enabling conditioning before generation. This paradigm is adopted by EditAR [mu2025editar] and VARSR [qu2025varsr]. While simple and model-agnostic, such strategies significantly increase sequence length and attention cost, limiting the model’s capacity for precise and instruction-aware editing. Furthermore, directly concatenating raw control tokens (e.g., in ControlVAR [controlvar]) may disrupt the pretrained generation behavior of the AR backbone. The reliance on semantically sparse VQ tokens further limits the effectiveness of prefilling-based approaches [harmon, blip3, unitok]. In this paper, we adopt prefilling-stage conditioning and propose a compact semantic method for better guidance. Our method applies to both next-token and next-set AR models with minimal architectural changes.
@@ 3.1 Preliminary
Autoregressive (AR) models extend the success of large language models [touvron2023llama, achiam2023gpt] to image generation and have become a popular approach in vision. These models convert images into discrete tokens using vector quantization (e.g., VQVAE [van2017neural]) and generate outputs by predicting tokens conditioned on preceding context. Recent AR approaches can be broadly categorized into two types: next-token prediction, which autoregressively generates individual tokens along a raster-scan sequence [llamagen, aim], and next-set prediction, which generates a set of tokens following specialized principles (e.g., VAR [var] predicts next-scale tokens in a coarse-to-fine manner). Despite the difference in basic token units (single token vs. a set of tokens), both follow a unified autoregressive formulation:
@@ 3.2 Compressed Semantic Prefilling
Our primary objective in the prefilling stage is to establish a conditioning prefix that is simultaneously computationally efficient (via short sequence length) and with rich semantics. To this end, we leverage powerful pre-trained Vision Foundation Models (VFM) $\mathcal{E}(\cdot)$ to encode the source condition image, and adopt DINOv2 [dinov2] in practice. However, this choice introduces a direct conflict with our goal of efficiency. For a standard $512\times 512$ input, they produce a sequence of 1024 tokens. As illustrated in Figure 1, this verbose prefix without compression leads to severe computational and memory overhead. To retain the rich semantics of VFMs while achieving a compact prefix, we propose a novel learnable token compression module.
@@
To ensure $\mathcal{P}_{k}(\cdot)$ learn to preserve critical semantic information during this compression, we train it with a lightweight upsampling module $\mathcal{U}_{k}$ implemented as a pixel shuffle operation [shi2016real, tokenshuffle] with a semantic preservation loss:
@@ 3.3 Semantic Alignment Guidance
Table 1: Quantitative results of C2I controllable generation on ImageNet [deng2009imagenet]. Values marked with ~ are estimated from histograms in the corresponding paper. Cells highlighted with $x$ and $x$ denote the best performance under the next-set and next-token autoregressive settings, respectively. SCAR-Uni denotes a unified model trained jointly across all conditions, rather than SCAR for each condition.
@@
Diff.
T2IAdapter [mou2024t2i] SD1.5 ~10.2 - ~9.9 - ~9.5 - ~9.3 - ~16.2 -
ControlNet [zhang2023adding] SD1.5 ~11.6 - ~9.2 - ~8.9 - ~8.6 - ~15.3 -
AR
ControlAR [li2024controlar] AiM-L 9.66 30.36 7.39 35.01 - - -
@@
Figure 3: Visualization of C2I controllable generation. Our SCAR demonstrates results respectively based on VAR [var] and LlamaGen [llamagen].
@@ 4.1 Experimental Setup
Dataset. We evaluate our method on controllable generation, covering class-to-image (C2I), text-to-image (T2I), and instruction editing tasks. For C2I controllable generation, models are trained on ImageNet-256 [deng2009imagenet] with five types of control conditions: Canny, Depth, Normal, HED, and Sketch. For T2I controllable generation, we use the MultiGen-20M [unicontrol] for training, considering Depth, Canny, HED, and Lineart as control conditions. For instruction editing, training is conducted on SEED-Edit-Unsplash [ge2024seed].
Training Details. (1) We use DINOv2-B [dinov2] as the image encoder due to its strong visual representations and scalability. (2) For C2I controllable generation at a resolution of $256\times 256$, we adopt two types of models: next-set VAR [var] and next-token LlamaGen [llamagen]. In this task, SCAR-Uni denotes a variant trained with uniformly sampled control types, supporting any control input at inference. (3) We use LlamaGen-XL with a T5 encoder [t5] at $512\times 512$ for both T2I controllable generation and instruction editing, using captions and editing instructions as text inputs, respectively. (4) We train VAR for 10 epochs and LlamaGen for 20 epochs on C2I controllable generation, 4 epochs on T2I controllable generation, and 2 epochs on instruction editing. (5) All models are trained on 8 NVIDIA H20 GPUs.
Table 2: Image generation quality and conditional consistency of T2I controllable generation on MultiGen-20M [unicontrol]. Our SCAR is based on LlamaGen-XL. Bold indicates the best result and underline indicates the second best. The results are conducted on 512 × 512 resolution.
@@
Diff.
GLIGEN [li2023gligen] 18.36 38.83 - 18.89 26.94 -
T2I-Adapter [mou2024t2i] 22.52 48.40 - 15.96 23.65 -
ControlNet [zhang2023adding] 17.76 35.90 15.41 76.21 14.73 34.65 17.44 70.54
Uni-ControlNet [zhao2023uni] 20.27 40.65 17.08 69.10 17.14 27.32 -
UniControl [qin2023unicontrol] 18.66 39.18 15.99 79.69 19.94 30.82 -
ControlNet++ [li2024controlnet++] 16.66 28.32 15.01 80.97 18.23 37.04 13.88 83.99
AR
ControlAR [li2024controlar] 14.61 29.01 10.53 85.63 17.51 37.08 12.41 79.22
EditAR [mu2025editar] 15.97 34.93 - 13.91 - -
@@
Evaluation and Metrics. For controllable generation, we evaluate models based on two main aspects: conditional consistency and image generation quality. Conditional consistency is measured by comparing the input condition image with the condition image extracted from the generated output. We use F1-Score for Canny and Sketch, Root Mean Square Error (RMSE) for Normal and Depth, and Structural Similarity Index Measure (SSIM) for HED and Lineart. Image generation quality is assessed using Fréchet Inception Distance (FID) [heusel2017gans]. Class-to-image (C2I) evaluations are conducted on the full 5050K validation set of ImageNet-256, while text-to-image (T2I) evaluations are performed on the official 55K validation set of MultiGen-20M [unicontrol].
For instruction editing, we adopt the PIE-Bench [ju2023direct], containing 700 samples covering 10 editing types. SCAR takes the source image and editing instructions as input to predict the target image, and follows the evaluation protocol in [ju2023direct]. Evaluation metrics cover three aspects: structure consistency, background preservation, and CLIP-based image-text alignment.

### v2

@@ Abstract
Recently, autoregressive (AR) models have shown strong potential in image generation, offering better scalability and easier integration with unified multi-modal systems compared to diffusion-based methods. However, extending AR models to general image editing remains challenging due to weak and inefficient conditioning, often leading to poor instruction adherence and visual artifacts. To address this, we propose SCAR, a Semantic-Context-driven method for AutoregRessive models. SCAR introduces two key components: Compressed Semantic Prefilling, which encodes high-level semantics into a compact and efficient prefix, and Semantic Alignment Guidance, which aligns the last visual hidden states with target semantics during autoregressive decoding to enhance instruction fidelity. Unlike decoding-stage injection methods, SCAR builds upon the flexibility and generality of vector-quantized-based prefilling while overcoming its semantic limitations and high cost. It generalizes across both next-token and next-set AR paradigms with minimal architectural changes. SCAR achieves superior visual fidelity and semantic alignment on both instruction editing and controllable generation benchmarks, outperforming prior AR-based methods while maintaining controllability. Code will be released at https://github.com/AMAP-ML/SCAR.
@@ 1 Introduction
The advent of large-scale generative models [56, 24, 16, 15, 36] has revolutionized the field of image editing [46, 89], opening up unprecedented possibilities for creative expression and content manipulation. Within this rapidly evolving landscape, two dominant generative paradigms have emerged: diffusion [19, 53] and autoregressive (AR) models [69, 71]. Diffusion models, renowned for their powerful text-to-image synthesis capabilities, demonstrate a strong grasp of semantic concepts, allowing for flexible and diverse edits based on user prompts. Concurrently, autoregressive models, owing to their excellent scaling properties and immense potential within unified multimodal generation and understanding model (UMM) architectures [35, 96], have achieved generation quality approaching that of top-tier diffusion models, establishing them as a critical and highly promising direction for future research.
@@
However, translating this raw generative power into versatile, general image editing capabilities remains a significant challenge for AR models. Compared to diffusion-based editing methods [4, 89, 27, 58], recent AR-based approaches [32, 50] have yet to achieve the same level of semantic control and general editing scope. These works can be broadly categorized into two types based on their conditioning strategies: (1) Decoding-stage injection (Figure 1.a), as in ControlAR [32], injects condition signals into the model’s intermediate layers. It performs excellent controllable generation, but the injected strong spatial guidance disrupts the autoregressive process for general editing (see Figure 6). (2) Prefilling-stage conditioning (Figure 1.b), employed by works like EditAR [50] and most UMMs  [81, 96, 25, 5], prepends visual tokens from the condition image to the input sequence. However, this method introduces debilitating computational overhead: prepending vector-quantized (VQ) tokens of the condition image can double the sequence length, increasing the computational cost, as shown in Figure 1.c. Moreover, VQ tokens are widely recognized as sparse-semantic [80, 6, 40], lacking the high-level representation needed for complex edits.
@@
First, we propose Compressed Semantic Prefilling, which replaces the inefficient VQ token prefix with compact, rich-semantic vision features extracted from the source image via a frozen vision foundation model (VFM) [52, 61, 10]. We introduce a learnable semantic compression module that drastically reduces the prefilling sequence length (e.g., $4\times$ compression from 1024 to 256 tokens) while preserving essential high-level semantics for editing. This significantly enhances computational efficiency and provides the model with a compact semantic understanding of the source content before generation, as shown in Figure 1.d. Notably, our semantic prefix remains robust even when compressed, while VQ token prefix suffers sharp performance drops under similar settings.
Second, to bridge the gap between sparse textual instructions and the desired dense visual guidance, we introduce a novel Semantic Alignment Guidance. Instead of relying solely on the sparse text condition, we use the visual features of the target edited image from the VFM as a dense semantic guide. Unlike the prior method [50] that distills supervision onto output VQ tokens, we apply an auxiliary constraint objective that aligns the autoregressive model’s last visual hidden states with the target’s semantic representation, providing a formulation more consistent with the causal decoding process. This provides a dense, in-context learning signal that steers the model’s internal reasoning toward the edit target.
@@ 2.1 Image Generation
Two dominant paradigms have emerged in image generation: diffusion models and AR models. Diffusion models synthesize images by iteratively denoising Gaussian noise, with DDPM [20] marking a major breakthrough. Subsequent work improves generation quality and efficiency [51, 64, 76, 91, 75, 59] through advances in sampling strategies and latent-space modeling. They have become the backbone of text-to-image and text-to-video generation [65, 67, 41, 24, 92, 77, 87, 7], typically using U-Net for denoising and CLIP [61, 78] or T5 [62] for text conditioning via cross-attention. Recent models like DiT [53] replace U-Net with Transformer backbones, achieving strong performance. Despite their success, diffusion models remain computationally expensive, motivating exploration of more efficient alternatives such as autoregressive approaches.
@@
In contrast, autoregressive (AR) models formulate image synthesis as a sequence modeling problem, predicting visual tokens step by step. Early works [73] focused on pixel-level generation. Building on the success of large language models [72, 1], recent approaches adopt discrete quantizers such as VQ-VAE [74] and VQ-GAN [11] to convert image patches into token indices, enabling next-token prediction over visual sequences. Recent advances have expanded this paradigm in two directions: (1) conventional next-token AR models [39, 43, 42], such as LlamaGen [69] and AiM [26], leverage large-scale transformer backbones for high-quality image synthesis; and (2) Visual Autoregressive Modeling (VAR), which reformulates generation as next-set prediction rather than raster-scan token prediction, offering improved scalability and global coherence [71, 44, 70, 15, 96]. Overall, AR-based approaches have achieved image quality comparable to diffusion models while being more efficient in sampling and deployment.
@@ 2.2 General Image Editing
General Image editing, including controllable and instruction editing, is a key focus in generative modeling, guided by external signals (e.g., text, edge maps, reference images). While early approaches incorporated class labels or attributes in GANs [47, 68] and VAEs [22, 88], recent diffusion-based methods [89, 49, 58] achieve fine-grained control via cross-attention or adapter modules. Building on this, text-guided instruction editing methods [17, 46, 4] further extend controllable generation to image manipulation. However, these approaches often involve costly inversion, require modality-specific tuning, or suffer from limited generalization across tasks and conditions.
In contrast, image editing with autoregressive (AR) models, including next-token and next-set paradigms, remains relatively underexplored. Recent AR-based methods can be broadly categorized by the stage at which control signals are injected: Decoding-stage injection [85, 32, 83] introduces control signals dynamically during generation, typically via cross-attention or feature modulation. These approaches offer strong controllability for pixel-level generation, but their complex architectural modifications often limit generalization to instruction-driven editing. Prefilling-stage conditioning [35, 81, 96, 50, 60, 30] prepends visual tokens from the source image to the input sequence, enabling conditioning before generation. This paradigm is adopted by EditAR [50] and VARSR [60]. While simple and model-agnostic, such strategies significantly increase sequence length and attention cost, limiting the model’s capacity for precise and instruction-aware editing. Furthermore, directly concatenating raw control tokens (e.g., in ControlVAR [30]) may disrupt the pretrained generation behavior of the AR backbone. The reliance on semantically sparse VQ tokens further limits the effectiveness of prefilling-based approaches [80, 6, 40]. In this paper, we adopt prefilling-stage conditioning and propose a compact semantic method for better guidance. Our method applies to both next-token and next-set AR models with minimal architectural changes.
@@ 3.1 Preliminary
Autoregressive (AR) models extend the success of large language models [72, 1] to image generation and have become a popular approach in vision. These models convert images into discrete tokens using vector quantization (e.g., VQVAE [74]) and generate outputs by predicting tokens conditioned on preceding context. Recent AR approaches can be broadly categorized into two types: next-token prediction, which autoregressively generates individual tokens along a raster-scan sequence [69, 26], and next-set prediction, which generates a set of tokens following specialized principles (e.g., VAR [71] predicts next-scale tokens in a coarse-to-fine manner). Despite the difference in basic token units (single token vs. a set of tokens), both follow a unified autoregressive formulation:
@@ 3.2 Compressed Semantic Prefilling
Our primary objective in the prefilling stage is to establish a conditioning prefix that is simultaneously computationally efficient (via short sequence length) and with rich semantics. To this end, we leverage powerful pre-trained Vision Foundation Models (VFM) $\mathcal{E}(\cdot)$ to encode the source condition image, and adopt DINOv2 [52] in practice. However, this choice introduces a direct conflict with our goal of efficiency. For a standard $512\times 512$ input, they produce a sequence of 1024 tokens. As illustrated in Figure 1, this verbose prefix without compression leads to severe computational and memory overhead. To retain the rich semantics of VFMs while achieving a compact prefix, we propose a novel learnable token compression module.
@@
To ensure $\mathcal{P}_{k}(\cdot)$ learn to preserve critical semantic information during this compression, we train it with a lightweight upsampling module $\mathcal{U}_{k}$ implemented as a pixel shuffle operation [66, 45] with a semantic preservation loss:
@@ 3.3 Semantic Alignment Guidance
Table 1: Quantitative results of C2I controllable generation on ImageNet [9]. Values marked with ~ are estimated from histograms in the corresponding paper. Cells highlighted with $x$ and $x$ denote the best performance under the next-set and next-token autoregressive settings, respectively. SCAR-Uni denotes a unified model trained jointly across all conditions, rather than SCAR for each condition.
@@
Diff. T2IAdapter [48] SD1.5 ~10.2 - ~9.9 - ~9.5 - ~9.3 - ~16.2 -
ControlNet [90] SD1.5 ~11.6 - ~9.2 - ~8.9 - ~8.6 - ~15.3 -
AR ControlAR [32] AiM-L 9.66 30.36 7.39 35.01 - - -
@@
Figure 3: Visualization of C2I controllable generation. Our SCAR demonstrates results respectively based on VAR [71] and LlamaGen [69].
@@ 4.1 Experimental Setup
Dataset. We evaluate our method on controllable generation, covering class-to-image (C2I), text-to-image (T2I), and instruction editing tasks. For C2I controllable generation, models are trained on ImageNet-256 [9] with five types of control conditions: Canny, Depth, Normal, HED, and Sketch. For T2I controllable generation, we use the MultiGen-20M [58] for training, considering Depth, Canny, HED, and Lineart as control conditions. For instruction editing, training is conducted on SEED-Edit-Unsplash [13].
Training Details. (1) We use DINOv2-B [52] as the image encoder due to its strong visual representations and scalability. (2) For C2I controllable generation at a resolution of $256\times 256$, we adopt two types of models: next-set VAR [71] and next-token LlamaGen [69]. In this task, SCAR-Uni denotes a variant trained with uniformly sampled control types, supporting any control input at inference. (3) We use LlamaGen-XL with a T5 encoder [62] at $512\times 512$ for both T2I controllable generation and instruction editing, using captions and editing instructions as text inputs, respectively. (4) We train VAR for 10 epochs and LlamaGen for 20 epochs on C2I controllable generation, 4 epochs on T2I controllable generation, and 2 epochs on instruction editing. (5) All models are trained on 8 NVIDIA H20 GPUs.
Table 2: Image generation quality and conditional consistency of T2I controllable generation on MultiGen-20M [58]. Our SCAR is based on LlamaGen-XL. Bold indicates the best result and underline indicates the second best. The results are conducted on 512 × 512 resolution.
@@
Diff. GLIGEN [31] 18.36 38.83 - 18.89 26.94 -
T2I-Adapter [48] 22.52 48.40 - 15.96 23.65 -
ControlNet [90] 17.76 35.90 15.41 76.21 14.73 34.65 17.44 70.54
Uni-ControlNet [93] 20.27 40.65 17.08 69.10 17.14 27.32 -
UniControl [57] 18.66 39.18 15.99 79.69 19.94 30.82 -
ControlNet++ [28] 16.66 28.32 15.01 80.97 18.23 37.04 13.88 83.99
AR ControlAR [32] 14.61 29.01 10.53 85.63 17.51 37.08 12.41 79.22
EditAR [50] 15.97 34.93 - 13.91 - -
@@
Evaluation and Metrics. For controllable generation, we evaluate models based on two main aspects: conditional consistency and image generation quality. Conditional consistency is measured by comparing the input condition image with the condition image extracted from the generated output. We use F1-Score for Canny and Sketch, Root Mean Square Error (RMSE) for Normal and Depth, and Structural Similarity Index Measure (SSIM) for HED and Lineart. Image generation quality is assessed using Fréchet Inception Distance (FID) [18]. Class-to-image (C2I) evaluations are conducted on the full 5050K validation set of ImageNet-256, while text-to-image (T2I) evaluations are performed on the official 55K validation set of MultiGen-20M [58].
For instruction editing, we adopt the PIE-Bench [21], containing 700 samples covering 10 editing types. SCAR takes the source image and editing instructions as input to predict the target image, and follows the evaluation protocol in [21]. Evaluation metrics cover three aspects: structure consistency, background preservation, and CLIP-based image-text alignment.